# ChauBoxplot and AdaptiveBoxplot: two R packages for boxplot-based outlier detection

### v1

@@ Abstract
Tukey’s boxplot is widely used for outlier detection; however, its classic fixed-fence rule tends to flag an excessive number of outliers as the sample size grows. To address this limitation, we introduce two new R packages, ChauBoxplot and AdaptiveBoxplot, which implement more robust methods for outlier detection. We also provide practical guidance, drawn from simulation results, to help practitioners choose suitable boxplot methods and balance interpretability with statistical reliability.
Keywords: Box-and-whisker plot, Chauvenet’s criterion, Chauvenet-type boxplot, Fence coefficient, Outlier detection, Sample size
@@ 1 Introduction
The box-and-whisker plot, first introduced by ?, is a fundamental and widely used tool in exploratory data analysis. It provides a concise summary of key distributional features, including the median, quartiles, and range, and visually identifies potential outliers. Because of its non-parametric nature and ease of interpretation, the boxplot is commonly applied in fields such as biomedicine, economics, engineering, psychology, and social sciences, as well as in quality control and educational research. A central function of the boxplot is outlier detection, which is performed by defining fences at fixed distances from the first quartile ($Q_{1}$) and third quartile ($Q_{3}$), and flagging any observations outside these fences as potential outliers. Tukey’s original boxplot sets the lower and upper fences at
@@
where $\mathrm{IQR}=Q_{3}-Q_{1}$ is the interquartile range, and $k=1.5$ is the standard fence coefficient. Although this outlier detection rule is simple and widely used, it lacks a rigorous theoretical foundation (?, ?, ?, ?, ?). Specifically, applying a fixed value of $k$ disregards the sample size and distributional characteristics, often leading to an excessive number of false positive outlier flags in larger samples.
To address the limitation imposed by a fixed fence coefficient, ? recently introduced the Chauvenet-type boxplot, which provides exact control of the outside rate per observation. For normally distributed data, by combining Tukey’s boxplot with Chauvenet’s criterion (?), their new fence coefficient is given as
@@
where $\Phi^{-1}$ denotes the standard normal quantile function and $n$ is the size of the data. Moreover, ? reframed the boxplot-based outlier detection as a multiple testing problem, treating each flagged observation as a statistical test against the majority of the data. Their methods provide a precise control over the family-wise error rate, the per-family error rate, or the false discovery rate. These methodological advances are now available in statistical software, with two new R packages—ChauBoxplot and AdaptiveBoxplot—offering convenient tools for practical implementation. On the other hand, the increased variety of options introduces new challenges, such as selecting suitable methods, interpreting error rates appropriately, and balancing simplicity with statistical rigor.
In this short paper, we present a focused review on the two representative R packages for boxplot-based outlier detection. Section 2 introduces ChauBoxplot, which implements the Chauvenet-type boxplot to provide exact control of the outside rate per observation, outlining its statistical foundation and practical application. Section 3 introduces AdaptiveBoxplot, which applies adaptive multiple testing procedures to improve outlier detection thresholds and accuracy. Section 4 presents a comparative study of the reviewed boxplot methods across a range of sample sizes. Section 5 offers practical guidance and recommendations for selecting boxplot-based outlier detection techniques in modern data analysis.
@@ 2 The ChauBoxplot Package
The Chauvenet-type boxplot introduces a novel fence coefficient that, on average, excludes half an observation from normally distributed data regardless of the sample size. Instead of using Tukey’s fixed fence coefficient, it employs the sample-size-adjusted alternative in (1) according to Chauvenet’s criterion. This adjustment integrates seamlessly into the standard boxplot construction, offering enhanced statistical reliability without sacrificing visual intuition or workflow convenience. For further technical details and the theoretical foundation of the Chauvenet-type coefficient, readers may refer to ?.
The ChauBoxplot package provides two main user interfaces for creating the Chauvenet-type boxplot. The chau boxplot()function acts as a direct drop-in replacement for base R’s boxplot(), with the only difference being the default $k=1.5$ replaced by $k_{n}^{\rm Chau}=\Phi^{-1}(1-0.25/n)/1.35-0.5$. For users within the ggplot2 ecosystem, the geom chau boxplot() function can also be implemented similarly to geom boxplot(). It provides a dedicated grammar of graphics layer that supports grouped, faceted, and highly customizable plots, making it well suited for advanced data analysis. Specifically, the new boxplot can be generated using the following syntax:
Ψlibrary(ChauBoxplot)
Ψ# For base R:
Ψchau_boxplot(x)
Ψ# For the ggplot2 ecosystem:
Ψggplot(df, aes(y = x, x = group)) + geom_chau_boxplot()
This compatibility allows researchers to adopt the new boxplot method with minimal changes to their existing code or analysis workflows.
The ChauBoxplot package offers comprehensive documentation, including a reference manual, a README file, and a set of vignettes that demonstrate usage with both synthetic and real-world datasets. These resources guide users through fundamental functions as well as advanced features, such as comparing Chauvenet’s criterion with Tukey’s classic rule and highlighting its advantages for large samples or non-normal data. The package is available on CRAN for all major R platforms and, released under the GPL-3 license, is actively maintained by the original authors to ensure it remains aligned with theoretical developments and evolving user needs. By combining statistically sound, the sample-size-adjusted outlier detection with an intuitive interface and thorough support materials, ChauBoxplot empowers researchers and analysts to conduct modern, reproducible data analysis across a broad range of scientific disciplines.
@@ 3 The AdaptiveBoxplot Package
The unified framework by ? extends the methodological advances of ? by systematically reframing the outlier detection as a multiple hypothesis testing problem. In this perspective, Tukey’s classic boxplot corresponds to an unadjusted testing procedure with a fixed per-comparison error rate, while the Chauvenet-type boxplot proposed by ? effectively controls the per-family error rate (PFER). Building on this insight, ? developed a generalized pipeline to construct boxplots that control more advanced error metrics including the family-wise error rate (FWER) and the false discovery rate (FDR). This pipeline proceeds in four steps:
  1. 1.
@@
This approach allows for the creation of fences that are adaptive not only to the sample size but also to the distributional characteristics of the data. Control of the FWER ensures a conservative bound ideal for strict error avoidance, while control of the FDR offers a powerful alternative that scales effectively with data complexity.
The AdaptiveBoxplot package implements this pipeline through two primary functions. The holm boxplot() function utilizes Holm’s step-down procedure (?) to control the FWER. This function is particularly well-suited for applications requiring strict conservatism, where avoiding even a single false positive is a priority. In contrast, the bh boxplot() function applies the Benjamini-Hochberg procedure (?) to control the FDR. This method offers a more powerful approach for large-scale or complex datasets, as it adaptively relaxes the fences when strong evidence of outliers is present, thereby maintaining high sensitivity without succumbing to the inflation of false discoveries typical of fixed-threshold methods.
@@
Ψ# Controlling the FWER at level 0.05:
Ψholm_boxplot(data, alpha = 0.05)
Ψ# Controlling the FDR at level 0.05:
Ψbh_boxplot(data, alpha = 0.05)
@@
The AdaptiveBoxplot package is freely available on CRAN under the GPL-3 license, supported by a concise reference manual that illustrates the function usage. As an actively maintained project, it ensures that these advanced, multiple-testing-based boxplots remain accessible to the R community. By embedding rigorous statistical inference into a familiar graphical tool, AdaptiveBoxplot facilitates more reproducible and defensible outlier detection. For more details on the unified $p$-value pipeline, readers are referred to ?.
4 Illustrating Example
@@
To illustrate the practical differences among the boxplot-based outlier detection methods, we conducted a simulation study comparing four representative approaches: Tukey’s boxplot, the Chauvenet-type boxplot, the FWER (Holm) boxplot, and the FDR (BH) boxplot. For each setting, we generated a dataset of size $n$, consisting of n3n-3 independent observations from the standard normal distribution and three outliers sampled from $N(5,\,0.5^{2})$. The sample sizes considered were $n=50$, 500500, 50005000 and 50,00050{,}000, allowing us to assess the performance across both small and large datasets. The data generation and analysis were implemented as follows:
Ψlibrary(ChauBoxplot)
Ψlibrary(AdaptiveBoxplot)
Ψset.seed(1000)
Ψn <- c(50, 500, 5000, 50000)
Ψm <- 3
Ψoutliers <- rnorm(m, 5, 0.5)
Ψpar(mfrow = c(4,4))
Ψoldpar = par(mar = c(1, 3.5, 2, 0.5), mgp = c(2, 0.5, 0))
Ψfor (i in 1:length(n)){
Ψ  x <- c(rnorm(n[i]-m), outliers)
Ψ  boxplot(x, main = "Tukey’s boxplot", ylab=paste("n =", n[i]))
Ψ  chau_boxplot(x)
Ψ  holm_boxplot(x, alpha = 0.05)
Ψ  bh_boxplot(x, alpha = 0.05)
Ψ}
@@
Figure 1: Comparison of boxplot-based outlier detection methods for normal data with three contaminated outliers. Each panel displays the results for one method (Tukey, Chauvenet-type, FWER [Holm] with rate 0.05, or FDR [BH] with rate 0.05) at $n=50$, 500500, 50005000 and 50 00050\,000, respectively.
@@
Figure 1 offers a comparative overview of the four boxplot-based outlier detection methods across a range of sample sizes. Each row represents a different sample size, and each column depicts a distinct method, with detected outliers shown as points beyond the whiskers. The three artificially introduced outliers are clearly visible in most panels. As the sample size increases, the classical boxplot with $k=1.5$ tends to misclassify an increasing number of central observations as outliers, especially for n=5,000n=5{,}000 and 50,00050{,}000, where many non-outlying values are incorrectly flagged. In contrast, the Chauvenet-type, FWER (Holm), and FDR (BH) boxplots only flag very few observations in total, demonstrating their robust performance regardless of the sample size.
@@
These results underscore the advantages of sample-size-adjusted and error-controlled boxplots for reliable outlier detection in moderate to large datasets. In contrast, the classical Tukey approach can lead to misleading conclusions when applied to large samples. Our simulation highlights the importance of adaptive strategies, such as those implemented in the ChauBoxplot and AdaptiveBoxplot packages, for principled and reproducible outlier identification.

### v2

@@ Abstract
Tukey’s boxplot is widely used for outlier detection; however, its classic fixed-fence rule tends to flag an excessive number of outliers as the sample size grows. To address this, we introduce two new R packages, ChauBoxplot and AdaptiveBoxplot, which implement more robust and statistically principled outlier detection methods. We illustrate their advantages and practical implications through comprehensive simulation studies and a real-world analysis of provincial university admission rates from China’s National College Entrance Examination. Based on these findings, we provide practical guidance to help practitioners select appropriate boxplot methods, achieving a balance between interpretability and statistical reliability.
keywords:
Box-and-whisker plot; Chauvenet’s criterion; Chauvenet-type boxplot; fence coefficient; outlier detection; sample size
@@ 1 Introduction
The box-and-whisker plot, first introduced by Tukey, (1977), is a fundamental and widely used tool in exploratory data analysis. It provides a concise summary of key distributional features, including the median, quartiles, and range, and visually identifies potential outliers. Due to its non-parametric nature and ease of interpretation, the boxplot is commonly applied in fields such as biomedicine, economics, engineering, psychology, and social sciences, as well as in quality control and educational research. A central function of the boxplot is outlier detection, which is performed by defining fences at fixed distances from the first quartile ($Q_{1}$) and the third quartile ($Q_{3}$), and flagging any observations outside these fences as potential outliers. Tukey’s original boxplot sets the lower and upper fences at
@@
where $\mathrm{IQR}=Q_{3}-Q_{1}$ is the interquartile range, and $k=1.5$ is the standard fence coefficient. Although this outlier detection rule is simple and widely used, it lacks a rigorous theoretical foundation (Hoaglin et al.,, 1986; Hoaglin and Iglewicz,, 1987; Frigge et al.,, 1989; Sim et al.,, 2005; Hofmann et al.,, 2017). Specifically, applying a fixed value of $k$ disregards the sample size and distributional characteristics, often leading to an excessive number of false positive outlier flags in larger samples.
To address the limitation imposed by a fixed fence coefficient, Lin et al., (2026) recently introduced the Chauvenet-type boxplot, which provides exact control of the outside rate per observation. For normally distributed data, combining Tukey’s boxplot with Chauvenet’s criterion (Chauvenet,, 1863) yields a new fence coefficient, given by
@@
where $\Phi^{-1}$ denotes the standard normal quantile function and $n$ is the size of the data. Moreover, Gang et al., (2026) reframed boxplot-based outlier detection as a multiple testing problem, treating each flagged observation as a statistical test against the majority of the data. Their methods provide precise control over the family-wise error rate, the per-family error rate, or the false discovery rate. These methodological advances are now available in statistical software, with two new R packages—ChauBoxplot and AdaptiveBoxplot—offering convenient tools for practical implementation. However, the increased variety of options introduces new challenges, such as selecting suitable methods, interpreting error rates appropriately, and balancing simplicity with statistical rigor.
In this short paper, we present a focused review of two representative R packages used for boxplot-based outlier detection. Section 2 introduces ChauBoxplot, which implements the Chauvenet-type boxplot to provide precise control over the outside rate for each observation and discusses its statistical foundation and practical applications. Section 3 presents AdaptiveBoxplot, which utilizes adaptive multiple testing procedures to enhance the accuracy and optimizes threshold selection in outlier detection. Section 4 presents comprehensive simulation studies comparing classical and modern boxplot approaches across a range of sample sizes. Section 5 presents a real-world application by analyzing provincial university admission rates from China’s National College Entrance Examination to illustrate the implications of boxplot-based outlier detection in practice. Finally, Section 6 offers practical guidance on selecting boxplot-based outlier detection techniques, summarizing the strengths and limitations of each method.
@@ 2 The ChauBoxplot package
The Chauvenet-type boxplot introduces a novel fence coefficient that, on average, excludes half an observation from normally distributed data regardless of the sample size. Instead of using Tukey’s fixed fence coefficient, it employs the sample-size-adjusted alternative in Equation (1) according to Chauvenet’s criterion. This adjustment integrates seamlessly into the standard boxplot construction, offering enhanced statistical reliability without sacrificing visual intuition or workflow convenience. For further technical details and the theoretical foundation of the Chauvenet-type coefficient, readers may refer to Lin et al., (2026).
The ChauBoxplot package provides two main user interfaces for creating the Chauvenet-type boxplot. The chau boxplot() function acts as a direct drop-in replacement for base R’s boxplot(), with the only difference being the default $k=1.5$ replaced by $k_{n}^{\rm Chau}=\Phi^{-1}(1-0.25/n)/1.35-0.5$. For users within the ggplot2 ecosystem, the geom chau boxplot() function can be used similarly to geom boxplot(). It provides a dedicated grammar of graphics layer that supports grouped, faceted, and highly customizable plots, making it well-suited for advanced data analysis. Specifically, the new boxplot can be generated using the following syntax:
ΨΨlibrary(ChauBoxplot)
ΨΨ# For base R:
ΨΨchau_boxplot()
ΨΨ# For the ggplot2 ecosystem:
ΨΨggplot() + geom_chau_boxplot()
This compatibility enables researchers to adopt the new boxplot method with minimal changes to their existing code or analysis workflows.
The ChauBoxplot package offers comprehensive documentation, including a reference manual, a README file, and a set of vignettes that demonstrate usage with both synthetic and real-world datasets. These resources include worked examples from our simulation studies and the analysis of provincial university admission rates presented in Sections 4 and 5, guiding users through both fundamental functions and advanced features. Users can explore comparisons between Chauvenet’s criterion and Tukey’s classic rule, and learn how ChauBoxplot provides advantages in handling large samples or non-normal data. The package is available on CRAN for all major R platforms under the GPL-3 license, and is actively maintained by the original authors to ensure it remains aligned with theoretical developments and evolving user needs. By combining statistically sound, sample-size-adjusted outlier detection with an intuitive interface and thorough support materials, ChauBoxplot empowers researchers and analysts to conduct modern, reproducible data analysis across a broad range of scientific disciplines.
@@ 3 The AdaptiveBoxplot package
The unified framework by Gang et al., (2026) extends the methodological advances of Lin et al., (2026) by systematically reframing outlier detection as a multiple hypothesis testing problem. From this perspective, Tukey’s classic boxplot corresponds to an unadjusted testing procedure with a fixed per-comparison error rate, while the Chauvenet-type boxplot proposed by Lin et al., (2026) effectively controls the per-family error rate (PFER). Building on this insight, Gang et al., (2026) developed a generalized pipeline to construct boxplots that control more advanced error metrics, including the family-wise error rate (FWER) and the false discovery rate (FDR). This pipeline proceeds in four steps:
  1. 1)
@@
This approach allows for the creation of fences that are adaptive not only to the sample size but also to the distributional characteristics of the data. Control of the FWER ensures a conservative bound that is ideal for strict error avoidance, while control of the FDR offers a powerful alternative that scales effectively with data complexity.
The AdaptiveBoxplot package implements this pipeline through two primary functions. The holm boxplot() function utilizes Holm’s step-down procedure (Holm,, 1979) to control the FWER. This function is particularly well-suited for applications requiring strict conservatism, where avoiding even a single false positive is a priority. In contrast, the bh boxplot() function applies the Benjamini-Hochberg procedure (Benjamini and Hochberg,, 1995) to control the FDR. This method offers a more powerful approach for large-scale or complex datasets, as it adaptively relaxes the fences when strong evidence of outliers is present, thereby maintaining high sensitivity without succumbing to the inflation of false discoveries typical of fixed-threshold methods.
@@
ΨΨ# Controlling the FWER at level 0.05:
ΨΨholm_boxplot(data, alpha = 0.05)
ΨΨ# Controlling the FDR at level 0.05:
ΨΨbh_boxplot(data, alpha = 0.05)
@@
The AdaptiveBoxplot package is freely available on CRAN under the GPL-3 license, supported by a concise reference manual that illustrates the function usage. As an actively maintained project, it ensures that these advanced, multiple-testing-based boxplots remain accessible to the R community. By embedding rigorous statistical inference into a familiar graphical tool, AdaptiveBoxplot facilitates more reproducible and defensible outlier detection. The package’s methods are demonstrated in the simulation study presented in Section 4, highlighting their practical advantages in controlling error rates for boxplot-based outlier detection. For more details on the unified $p$-value pipeline, readers may refer to Gang et al., (2026).
4 Simulated data example
@@
To illustrate the practical differences among the boxplot-based outlier detection methods, we conducted a simulation study comparing four representative approaches: Tukey’s boxplot, the Chauvenet-type boxplot, the FWER (Holm) boxplot, and the FDR (BH) boxplot. For each setting, we generated a dataset of size $n$, consisting of n3n-3 independent observations from the standard normal distribution $\mathcal{N}(0,1)$ and three outliers sampled from $\mathcal{N}(5,0.5^{2})$. The sample sizes considered were $n=50$, 500500, 50005000 and $50000$, respectively, allowing us to assess the performance across both small and large datasets. The data generation and analysis were implemented as follows:
ΨΨlibrary(ChauBoxplot)
ΨΨlibrary(AdaptiveBoxplot)
ΨΨn <- c(50, 500, 5000, 50000)
ΨΨm <- 3
ΨΨoutliers <- rnorm(m, 5, 0.5)
ΨΨpar(mfrow = c(4,4))
ΨΨoldpar = par(mar = c(1, 3.5, 2, 0.5), mgp = c(2, 0.5, 0))
ΨΨfor (i in 1:length(n)){
Ψ  Ψx <- c(rnorm(n[i]-m), outliers)
Ψ  Ψboxplot(x, main = "Tukey’s boxplot", ylab=paste("n =", n[i]))
Ψ  Ψchau_boxplot(x)
Ψ  Ψholm_boxplot(x, alpha = 0.05)
Ψ  Ψbh_boxplot(x, alpha = 0.05)
ΨΨ}
@@
Figure 1: Comparison of boxplot-based outlier detection methods for normal data with three contaminated outliers. Each panel displays the results for one method (Tukey, Chauvenet-type, FWER [Holm] with rate 0.05, or FDR [BH] with rate 0.05) at $n=50$, 500500, 50005000 and $50000$, respectively.
@@
Figure 1 offers a comparative overview of the four boxplot-based outlier detection methods across a range of sample sizes. Each row represents a different sample size, and each column depicts a distinct method, with detected outliers shown as points beyond the whiskers. The three artificially introduced outliers are clearly visible in most panels. As the sample size increases, the classical boxplot with $k=1.5$ tends to misclassify an increasing number of central observations as outliers, especially for $n=5000$ and $50000$, where many non-outlying values are incorrectly flagged. In contrast, the Chauvenet-type, FWER (Holm), and FDR (BH) boxplots only flag very few observations in total, demonstrating their robust performance regardless of the sample size. Finally, we note that Figure 1 is fully reproducible: by setting set.seed(1000) for random number generation, the provided R code can be run directly to obtain the figure as shown.
@@
These results underscore the advantages of sample-size-adjusted and error-controlled boxplots for reliable outlier detection in moderate to large datasets. In contrast, the classical Tukey approach can lead to misleading conclusions when applied to large samples. Our simulation study demonstrates that the adaptive strategies implemented in the ChauBoxplot and AdaptiveBoxplot packages offer principled control of error rates and maintain robust outlier identification across a wide range of sample sizes. These findings highlight the practical value of using statistically grounded methods for reproducible and defensible outlier detection.
5 Real data example
To further illustrate the application of boxplot-based outlier detection, we analyzed provincial admission rates from the 2024 National College Entrance Examination (Gaokao) in China. The Gaokao is among the world’s largest standardized tests, with a record 13.53 million participants in 2024. In this analysis, we focused on admission rates to two groups of elite universities: the “Project 985”, which comprises 39 institutions, and the “Project 211”, which encompasses 115 universities. These measures are widely used as indicators of regional educational opportunity. Table 1 presents the number of examinees and admission rates by province, ordered by the Project 985 admission rate. Municipalities such as Tianjin, Beijing, and Shanghai have the highest Project 985 rates, reflecting concentrated resources, while provinces such as Jiangxi, Guizhou, and Anhui have among the lowest, underscoring regional disparities.
Table 1: Provincial summary of Gaokao 2024 examinee counts, admissions to the Project 985 and Project 211 universities, and associated admission rates. The table is sorted in descending order by the Project 985 admission rate. Admission rates denote the percentage of total examinees in each province admitted to the respective university category.
 Province Total Examinees 985 Admissions     985 Rate 211 Admissions     211 Rate
Tianjin 74000 4410 6.0% 9588 13.0%
Beijing 68000 3604 5.3% 11319 16.7%
Shanghai 58000 2541 4.4% 6520 11.2%
Jilin 132000 4230 3.2% 10575 8.0%
Qinghai 59000 1579 2.7% 6159 10.4%
Ningxia 82000 1686 2.1% 6306 7.7%
Shaanxi 350000 6983 2.0% 18824 5.4%
Liaoning 203000 4129 2.0% 10593 5.2%
Zhejiang 405000 7106 1.8% 15905 3.9%
Chongqing 350000 6198 1.8% 15939 4.6%
Hubei 520000 9513 1.8% 23782 4.6%
Shanxi 354000 6019 1.7% 18690 5.3%
Shandong 1000000 16544 1.7% 35344 3.5%
Fujian 247000 4098 1.7% 11066 4.5%
Jiangsu 480000 7633 1.6% 26333 5.5%
Heilongjiang 211000 3422 1.6% 11462 5.4%
Xinjiang 230000 3516 1.5% 16130 7.0%
Hainan 73000 1125 1.5% 3849 5.3%
Guangxi 480000 6768 1.4% 25944 5.4%
Guangdong 760000 10558 1.4% 23096 3.0%
Henan 1360000 18471 1.4% 54182 4.0%
Gansu 260000 3483 1.3% 8126 3.1%
Sichuan 830000 10857 1.3% 31847 3.8%
Xizang 39000 511 1.3% 3850 9.9%
Hunan 700000 9236 1.3% 23397 3.3%
Yunnan 406000 5040 1.2% 16201 4.0%
Hebei 880000 10143 1.2% 32768 3.7%
Inner Mongolia 230000 2609 1.1% 10434 4.5%
Anhui 670000 7344 1.1% 26552 4.0%
Guizhou 506000 5392 1.1% 23365 4.6%
Jiangxi 650000 6430 1.0% 24647 3.8%
More specifically, we use the admission rates for Project 985 and Project 211 universities from https://www.gk100.com/read_12602336.htm for this analysis. All visualizations are produced with ggplot2, as Section 4 already covers base R graphics. Because AdaptiveBoxplot supports only base R, this example focuses on comparing Tukey’s classical boxplot with the sample-size-adjusted Chauvenet-type boxplot. The complete analysis code, including scripts to reproduce Figure 2 for both Project 985 and Project 211 admission rates, is available at https://github.com/tiejuntong/ChauBoxplot, ensuring the analysis is entirely reproducible.
ΨΨlibrary(ggplot2)
ΨΨlibrary(ChauBoxplot)
ΨΨprovince <- c("Tianjin", ..., "Jiangxi")
ΨΨrate.985 <- c(6.0, ..., 1.0)
ΨΨgaokao <- data.frame(province, rate.985, rate.211)
ΨΨtheme_box <- theme(legend.position = "none",
  ΨΨplot.margin = unit(c(0, 0, 0, 0), "inches"))
ΨΨggplot(gaokao, aes(x = "", y = rate.985)) +
$Ψ$  geom_boxplot(width = 0.5) +
$Ψ$  scale_x_discrete(breaks = NULL) + theme_box +
$Ψ$  labs(title = "Tukey’s Boxplot", subtitle = "985 Rate (%)")
ΨΨggplot(gaokao, aes(x = "", y = rate.985)) +
$Ψ$  geom_chau_boxplot(width = 0.5) +
$Ψ$  geom_text(data = subset(gaokao, province == "Qinghai"),
            mapping = aes(x = "", y = rate.985, label = province),
            nudge_x = 0.25, size = 4.5, color = "blue", vjust = 0.5) +
$Ψ$  scale_x_discrete(breaks = NULL) + theme_box +
$Ψ$  labs(title = "ChauBoxplot", subtitle = "985 Rate (%)")
Figure 2: Tukey’s boxplot and the Chauvenet-type boxplot of the Gaokao admission rates in 2024. The left two panels are for the Project 985 university admission rates, while the right two panels are for the Project 211 university admissions rates.
For the Project 985 admission rates, the first quartile ($Q_{1}$) is 1.3% and the third quartile ($Q_{3}$) is 1.9%, giving an interquartile range (IQR) of 0.6%. Tukey’s boxplot flags observations above the upper fence, calculated as $Q_{3}+1.5\times\mathrm{IQR}=2.8\%$, identifying Tianjin (6.0%), Beijing (5.3%), Shanghai (4.4%), and Jilin (3.2%) as outliers. In contrast, the Chauvenet-type boxplot uses a sample-size-adjusted coefficient, $k_{n}^{\rm Chau}={\Phi^{-1}(1-0.25/n)}/{1.35}-0.5=1.28$ with $n=31$, yielding an upper fence of $Q_{3}+1.28\times\mathrm{IQR}=2.67\%$. This approach flags the same high-rate provinces and, due to the slightly lower threshold, also identifies Qinghai (2.7%) as a borderline outlier. These results demonstrate that the adaptive Chauvenet-type boxplot is more sensitive to moderate but meaningful deviations, providing a nuanced view of regional disparities in educational opportunity.
For the Project 211 admission rates, both Tukey’s boxplot and the Chauvenet-type boxplot identified Beijing (16.7%), Tianjin (13.0%), Shanghai (11.2%), Qinghai (10.4%), and Xizang (9.9%) as outliers. These provinces are characterized by exceptionally high admission rates to the Project 211 universities, reflecting a strong concentration of educational resources in provincial-level municipalities as well as distinct regional patterns in Qinghai and Xizang. Overall, these results show that both traditional and adaptive boxplot methods consistently highlight provinces with significantly greater access to elite higher education, while the Chauvenet-type approach offers additional sensitivity to borderline cases.